OpenStudy (anonymous):

2x + y = 1 4x + 2y = −2 x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions. @nathan917 ?

OpenStudy (anonymous):

@amistre64 ?

OpenStudy (nathan917):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1......................(i) 4x + 2y = −1 ..................(ii) Since we have to find x, there y should be eliminated, multiply (i) by 2 4x +2y = 2.....................(iii) subtract (ii) from (iii) 0= 3 ,it cannot be possible Hence there is no solution, because both are parallel..........................Ans There is no x value as there is no solution.

OpenStudy (anonymous):

Thanks @nathan917 :)

OpenStudy (nathan917):

Np. Tanks for mentioning me.

OpenStudy (nathan917):

Thanks*